OpenStudy (anonymous):

f(x) is a continuous fxn. f(x)+f(2x+y)+5xy=f(3x-y)+2x^{2}+1 for all x belongs to R then which of these holds true?? (a) f is many one (b) f has no minima (c) f is neither even nor odd (d) f is bounded please help me with this problem

OpenStudy (anonymous):

\(\mathsf{f(x) = 1- \frac 1 2 x^2}\) satisfies the equation. But it's a even function, a parabola and has no defined Minima. I don't know what you mean by bounded, but I think it's probably not bounded. And because it's a parabola yes it's a many one function as well. A and B.

OpenStudy (anonymous):

yeah thanxx i got it....thanx a lot..;) u r really genious.:)

OpenStudy (anonymous):

You're welcome :-)

OpenStudy (anonymous):

but how this function came into ur mind...can u please tell??

OpenStudy (anonymous):

Umm first I defined y = 0. \[f(x) + (2x) - f(3x) = 2x^2 + 1\tag1\] I checked the function for x=0 \[f(0) = 1\] Then I just guessed a function, I really don't have much working to show. If f(0) =1, I presumed f(x) must be something like \(ax^1 + bx^2 +\ldots +nx^n + 1\). I used the equation (1) to get the idea the further idea of the function. I only needed to get one function that satisfies the equation \(f(x)+f(2x+y)+5xy=f(3x-y)+2x^{2}+1\), for that I simply assumed the function to be \(ax^2 +1\).